### v3

# Relations between average shortest path length and another centralities in graphs

Abstract
Relations between average shortest path length and radiality, closeness, stress centralities and average clustering coefficient were obtained for simple connected graphs.
keywords:
Networks, centralities, local and global properties of graphs, average shortest path length, Watts-Strogatz clustering coefficient.
1 Introduction.
Centrality measures are important characteristics in network science which contain intrinsic hidden information about networks. Centrality is a local (with relation to a vertex) or global (with relation to a whole graph) real functions on graphs. There exist many centralities [1][3]: local efficiency, radiality, closeness, betweenness, stress centralities, etc. One of the most important centralities for “real” networks (or small-world networks) are average clustering coefficient and average shortest path length [4]. The network is called small-world if it has small average shortest path length with relation to the size and big average clustering coefficient. This measures are separately calculated to prove that current network is small-world. In this article we proved that for geodetic graphs (or graphs with odd length cycles) with additional condition there is relation between average shortest path length and average clustering coefficient.
2 Main definitions.
All subsequent definitions are given for a simple, connected, undirected graph $G$. Let’s denote by
  • $V(G)$ the set of vertices,   $E(G)$ the set of edges and   $A=\{a_{ij}\}$ adjacency matrix of graph $G$.
  • Neighborhood $N(v)$ — the induced subgraph in $G$ on vertices which adjacent to the vertex $v$,
  • $N^{\prime}(v)$ induced subgraph in $G$ on vertices $=V\big{(}N(v)\big{)}\bigcup\{v\}$,
  • $\bar{f}(x_{1},x_{2},...,x_{k})$, where $f$ is any function $V\times V\times...\times V\rightarrow\mathbb{R}$, the restriction of this function on $N^{\prime}(v)$ (for example $\bar{L}(x,y)$ will be the average shortest path between $x$ and $y$ with restriction to subgraph $N^{\prime}(v)$),
  • $d_{i}=deg(v_{i})$,
  • $n=|V(G)|,\;\;m=|E(G)|$,
  • $X(i)=X(v_{i})$ for any $X$ — set or function corresponding to vertex $v_{i}$,
Let’s give definitions of centralities:
  1. 1.
    Diameter $diam(G)=\max_{s,t\in V(G)}dist(s,t)$.
  2. 2.
    Average shortest path length $L(G)=\frac{1}{n(n-1)}\sum\limits_{s\neq t}dist(s,t)$.
  3. 3.
    Local cluster coefficient
    $c_{i}=c(i)=\frac{\text{number of edges in }N(i)}{\text{maximum possible numberof edges in }N(i)}=\frac{2|E(N(i))|}{d_{i}(d_{i}-1)}$.
  4. 4.
    Average clustering coefficient
    $C_{WS}(G)=\frac{1}{n}\sum\limits_{i\in V(G)}c_{i}=\frac{1}{n}\sum\limits_{i\in V(G)}\frac{2|E(N(i)))|}{d_{i}(d_{i}-1)}=\frac{1}{n}\sum\limits_{i\in V(G)}\frac{\sum\limits_{j,k\in V(G)}a_{ij}a_{jk}a_{ki}}{d_{i}(d_{i}-1)}$.
  5. 5.
    Global clustering coefficient
    $C(G)=\frac{\text{number of closed triplets in $G$}}{\text{number of all triplets in $G$}}=\frac{\sum\limits_{i,j,k\in V(G)}a_{ij}a_{jk}a_{ki}}{\sum\limits_{i\in V(G)}d_{i}(d_{i}-1)}$.
  6. 6.
    Closeness centrality $Clo(v)=\frac{n-1}{\sum\limits_{t\in V(G)}dist(v,t)}$.
  7. 7.
    Radiality $Rad(v)=\frac{\sum\limits_{t\in V(G),t\neq v}(diam(G)+1-dist(v,t))}{n-1}$.
  8. 8.
    Stress $Str(i)=\sum\limits_{s,t\in V(G),\;s\neq t\neq i}\sigma_{st}(i)$, where $\sigma_{st}(i)$ is the total number of shortest paths from $s$ to $t$ which contains vertex $i$.
Note that all centralities are non-negative and $c_{i},C_{WS},\operatorname{Clo}(v)$ are less or equal 1.
Also let’s give the definition of geodetic graph:
Definition 1.
If there exists the unique shortest path between any two vertices in $G$ then the graph $G$ is called geodetic.
This definition is equivalent to the condition then there is no even cycles in a graph.
3 Main results.
Let’s consider a induced subgraph $G^{\prime}\subset G$. In general $G^{\prime}$ can be not connected graph. In this case let’s define the average shortest path length for vertices of $G^{\prime}$ with relation to the distance $\operatorname{dist}$ in the ambient graph $G$. Let’s call $L\big{(}N(i)\big{)}$ the local average shortest path length for the vertex $i$.
Let’s start with simple relations: let’s proof the relation between local shortest path length and local clustering coefficient:
Lemma 1.
$L(N(i))=2-c_{i}.$
{addmargin}
[1em]0em
Proof.
$L(N(i))=\frac{1}{d_{i}(d_{i}-1)}\sum\limits_{s,t\in N(i),s\neq t}dist(s,t)=\frac{1}{d_{i}(d_{i}-1)}\sum\limits_{(s,t)\in E(N(i))}dist(s,t)+\sum\limits_{s,t\in N(i),(s,t)\notin E(N(i))}dist(s,t)=$
$=\frac{1}{d_{i}(d_{i}-1)}(2|E(N(i))|+\sum\limits_{(s,i),(i,t)\in E(G),(s,t)\notin E(G)}dist(s,t))=$
$=\frac{1}{d_{i}(d_{i}-1)}(2|E(N(i))|+2(d_{i}(d_{i}-1)-2|E(N(i))|))=2-c_{i}.$
Note that shortest paths for vertices in $N(i)$ are defined corresponding to whole graph $G$. ∎
Averaging by $i$ we obtain simple corollary about the relation between local shortest path length and average clustering coefficient.
Corollary 1.
$C_{WS}(G)=2-\frac{1}{n}\sum\limits_{i\in V(G)}L\big{(}N(i)\big{)}.$
Let’s proof the relation between shortest path length in subgraph $N^{\prime}(i)$ and shortest path length in $N(i)$.
Lemma 2.
$L\big{(}N^{\prime}(i)\big{)}=\frac{(d_{i}-1)L\big{(}N(i)\big{)}+2}{d_{i}+1}.$
{addmargin}
[1em]0em
Proof.
By definition
$L\big{(}N^{\prime}(i)\big{)}=\frac{1}{(d_{i}+1)d_{i}}\sum\limits_{s,t\in V(N^{\prime}(i)),\;s\neq t}\operatorname{dist}(s,t)=\frac{d_{i}-1}{d_{i}+1}L\big{(}N(i)\big{)}+\frac{2}{d_{i}+1}.$
Let’s proof the relation between shortest path length in a induced subgraph and shortest path length in ambient graph if induced subgraph is obtained from ambient graph by deleting one vertex.
Theorem 1.
Let a graph $G$ is obtained from a connected simple graph $G^{\prime}$ by deleting one vertex and $|V(G)|=n$. Then
$L(G^{\prime})\geq\frac{n}{n+1}L(G),$
where the average shortest path length $L(G)$ is defined in the ambient graph $G^{\prime}$, if $G$ is not connected.
{addmargin}
[1em]0em
Proof.
Let’s define the deleted vertex by $v$. By the triangle inequality $\forall s,t\in V(G):\operatorname{dist}(s,v)+\operatorname{dist}(v,t)\geq\operatorname{dist}(s,t)$, where the equality holds if, there are no paths from $s$ to $t$ in $G$. Therefore,
$\sum\limits_{s,t\in V(G),s\neq t}\big{(}\operatorname{dist}(s,v)+\operatorname{dist}(v,t)\big{)}\geq\sum\limits_{s,t\in V(G),s\neq t}\operatorname{dist}(s,t)$
$\frac{2(n-1)}{n(n-1)}\sum\limits_{t\in V(G)}\operatorname{dist}(v,t)\geq\frac{1}{n(n-1)}\sum\limits_{s,t\in V(G),s\neq t}\operatorname{dist}(s,t)$
$\frac{2}{n}\sum\limits_{t\in V(G)}\operatorname{dist}(v,t)\geq L(G)$
Then,
$L(G^{\prime})=\frac{1}{(n+1)n}\sum\limits_{s,t\in V(G^{\prime}),s\neq t}\operatorname{dist}(s,t)=\frac{1}{(n+1)n}\Big{(}2\sum\limits_{t\in V(G)}\operatorname{dist}(v,t)+\sum\limits_{s,t\in V(G),s\neq t}\operatorname{dist}(s,t)\Big{)}=$
$=\frac{1}{n+1}\frac{2}{n}\sum\limits_{t\in V(G)}\operatorname{dist}(v,t)+\frac{n-1}{n+1}L(G)\geq\frac{n}{n+1}L(G)$
Note that if $G$ consists of $n$ isolated vertices then the equality holds. ∎
We obtain corollaries
Corollary 2.
Let a graph $G$ is obtained from a connected simple graph $G^{\prime}\subset H$ by deleting one vertex, $|V(G)|=n$ and $H$ is connected and simple. Then
$L(G^{\prime})\geq\frac{n}{n+1}L(G),$
where the average shortest path lengths $L(G)$ and $L(G^{\prime})$ are defined in the ambient graph $H$, if corresponded graphs are not connected.
{addmargin}
[1em]0em
Proof.
The proof is the same as for the previous theorem and also if $G$ consists of $n$ isolated vertices then the equality holds. ∎
Corollary 3.
$L\big{(}N^{\prime}(i)\big{)}\geq\frac{d_{i}}{d_{i}+1}L\big{(}N(i)\big{)}.$
Let’s proof the relation between shortest path length in a induced subgraph and shortest path length in ambient graph.
Theorem 2.
Let’s $G^{\prime}\subset G$ be induced subgraph and $|V(G)|=n,|V(G^{\prime})|=n+k$. Then
$L(G^{\prime})\geq\frac{n}{n+k}L(G),$
where the average shortest path length $L(G)$ is defined in the ambient graph $G^{\prime}$, if $G$ is not connected.
{addmargin}
[1em]0em
Proof.
Let’s construct the graph $G^{\prime}$ from $G$ by adding sequentially $k$ vertices and corresponded edges. Let’s first sequentially add vertices adjacent to vertices of the graph $G$. Adding these vertices one by one we obtain a sequence of graphs $G_{1},G_{2},...,G_{p},$ where $|V(G_{i})|=n+i$. Further, let’s add is the same way vertices adjacent to vertices of the graph $G_{p}$ and so on. In the end we obtain the graph $G^{\prime}$. By previous corollary
$L(G^{\prime})\geq\frac{n+k-1}{n+k}L(G_{k-1})\geq\frac{n+k-1}{n+k}\frac{n-k-2}{n-k-1}L(G_{k-2})=\frac{n-k-2}{n+k}L(G_{k-2})\geq\cdots\geq\frac{n}{n+k}L(G).$
Let’s proof the relation between average closeness centrality and average shortest path length:
Lemma 3.
$L(G)\geq\frac{n}{\sum\limits_{v\in V(G)}\operatorname{Clo}(v)}.$
{addmargin}
[1em]0em
Proof.
By the inequality of harmonic mean and arithmetic mean
$\frac{1}{n}\sum\limits_{v\in V(G)}Clo(v)=\frac{1}{n}\sum\limits_{v\in V(G)}\frac{n-1}{\sum\limits_{t\in V(G)}dist(v,t)}\geq\frac{n(n-1)}{\sum\limits_{v,t\in V(G)}dist(v,t)}=\frac{1}{L(G)}.$
Note that an equality holds when all average shortest path lengths from any vertex to all remaining vertices are equal. ∎
Let’s proof the relation between average shortest path length and average radiality:
Lemma 4.
$L(G)=\operatorname{diam}(G)+1-\frac{1}{n}\sum\limits_{v\in V(G)}{\operatorname{Rad}}(v).$
{addmargin}
[1em]0em
Proof.
The proof holds from definition
$\frac{1}{n}\sum\limits_{v\in V(G)}{Rad}(v)=\frac{1}{n}\sum\limits_{v\in V(G)}\frac{(n-1)(diam(G)+1)-\sum\limits_{t\in V(G),\;t\neq v}dist(v,t))}{n-1}=diam(G)+1-L(G).$
Now let’s prove theorem about a relation between average clustering coefficient and radiality using previous theorem.
Theorem 3.
$\frac{C_{WS}(G)}{2}\geq\frac{1}{n}\sum\limits_{i\in V(G)}\frac{\sum\limits_{v\in N(i)}\overline{\operatorname{Rad}}(v)}{d_{i}}-2+\frac{\#\{N(i)\text{ whichare complete graphs}\}}{n}.$
{addmargin}
[1em]0em
Proof.
By lemmas 3: $L(N^{\prime}(i))\geq\frac{d_{i}}{d_{i}+1}L(N(i))\geq\frac{1}{2}L(N(i))=1-\frac{1}{2}c_{i}$. Therefore,
$\frac{1}{d_{i}}\sum\limits_{v\in N(i)}\overline{\operatorname{Rad}}(v)=\operatorname{diam}(N^{\prime}(i))+1-L(N^{\prime}(i))\leq\operatorname{diam}(N^{\prime}(i))+\frac{1}{2}c_{i}=\frac{1}{2}c_{i}+2-\chi_{K_{d_{i}}}(N(i)),$
where $\chi_{K_{d_{i}}}(N(i))=\begin{cases}1&\text{if $N(i)=K_{d_{i}}$}\\ 0&\text{otherwise}\end{cases}$. Averaging by $i$ ends the proof. ∎
Let’s prove a theorem about a connection between the average stress centrality and average shortest path length for geodetic graphs.
Theorem 4.
If $G$ is geodetic, then
$L(G)=1+\frac{1}{n(n-1)}\sum\limits_{i\in V(G)}\operatorname{Str}(i).$
{addmargin}
[1em]0em
Proof.
Let’s define
$\chi_{st}(i)=\begin{cases}1&\text{if $i\neq s\neq t$ is the vertex of the shortest path between $s$ and $t$,}\\ 0&\text{otherwise.}\end{cases}$
Thus, $\operatorname{Str}(i)=\sum\limits_{s,t\in V(G)}\chi_{st}(i)$. If there exists the unique shortest path between any two vertices in $G$, then $\operatorname{dist}(s,t)=\sum\limits_{i\in V(G)}\chi_{st}(i)+1$ (otherwise, $\operatorname{dist}(s,t)\leq\sum\limits_{i\in V(G)}\chi_{st}(i)+1$). Therefore, for any $i$
$\sum\limits_{s,t\in V(G)}\operatorname{dist}(s,t)=2|E|+\sum\limits_{s,t\in V(G),\,\operatorname{dist}(s,t)\geq 2}\operatorname{dist}(s,t)=2|E|+\sum\limits_{s,t\in V(G),\,\operatorname{dist}(s,t)\geq 2}\Big{(}\sum\limits_{i\in V(G)}\chi_{st}(i)+1\Big{)}=$
$=2|E|+\sum\limits_{i\in V(G)}\operatorname{Str}(i)+n(n-1)-2|E|=\sum\limits_{i\in V(G)}\operatorname{Str}(i)+n(n-1).$
Corollary 4.
For any simple connected $G$
$L(G)\leq 1+\frac{1}{n(n-1)}\sum\limits_{i\in V(G)}\operatorname{Str}(i).$
First let’s proof theorem about a connection between average clustering coefficient and stress centrality and we will use it for a connection between the average shortest path length and the average local clustering coefficient further.
Theorem 5.
If there is no pendant vertices in graph, then
$C_{WS}(G)\geq 1-\frac{1}{n}\sum\limits_{i\in V(G)}\frac{Str(i)}{d_{i}(d_{i}-1)}.$
{addmargin}
[1em]0em
Proof.
Note that $\forall j,k\in N(i):(j,k)\notin E(N(i))$ the shortest path between $j$ and $k$ is $j\rightarrow i\rightarrow k$. Therefore,
$Str(i)\geq 2(\frac{d_{i}(d_{i}-1)}{2}-|E(N(i))|),$
$\frac{1}{d_{i}(d_{i}-1)}Str(i)\geq 1-c_{i},$
Averaging by $i$
$C_{WS}(G)\geq\frac{1}{n}\sum\limits_{i\in V(G)}(1-\frac{Str(i)}{d_{i}(d_{i}-1)}).$
Note that for $diam(G)=2$ holds an equality. ∎
Corollary 5.
In the general case
$C_{WS}(G)\geq 1-\frac{1}{n}\sum\limits_{i\in V(G)}\frac{Str(i)}{d_{i}(d_{i}-1)}-\frac{\text{number of pendant vertices}}{n}.$
{addmargin}
[1em]0em
Proof.
Let’s for pendant vertex define $\frac{Str(i)}{d_{i}(d_{i}-1)}$ as 0, then in the inequality $\frac{1}{d_{i}(d_{i}-1)}Str(i)\geq 1-c_{i}$ in the right side will be 1 and in the left side 0, thus if we add in the left side 1 for every pendant vertex, will be right equality. ∎
Now let’s prove a theorem about a connection between the average shortest path length and the average local clustering coefficient for geodetic graphs.
Theorem 6.
If $G$ is geodetic and $\forall i,j\in V(G)$ hold, if $d_{i}\leq d_{j}$ then $\operatorname{Str}(i)\leq\operatorname{Str}(j)$, then
$1-C_{WS}(G)\leq\frac{1}{n}\sum\limits_{i\in V(G)}\frac{\big{(}L(G)-1\big{)}(n-1)}{d_{i}(d_{i}-1)}+\frac{\text{number of pendant vertices}}{n}.$
{addmargin}
[1em]0em
Proof.
Let’s re-numerate vertices such that $\forall i\leq j:d_{i}\leq d_{j}$. Then for $i\leq j$ hold $\operatorname{Str}(i)\leq\operatorname{Str}(j)$ and $d_{i}(d_{i}-1)\leq d_{j}(d_{j}-1)$. By theorems 4 and 5 for the case if there is no pendant vertices in graph
$1-C_{WS}(G)\leq\frac{1}{n}\sum\limits_{i\in V(G)}\frac{Str(i)}{d_{i}(d_{i}-1)}\stackrel{{\scriptstyle\text{Chebyshev's sum inequality}}}{{\leq}}$
$\hskip 70.0pt\leq\Big{(}\frac{1}{n}\sum\limits_{i\in V(G)}{Str(i)}\Big{)}\Big{(}\frac{1}{n}\sum\limits_{i\in V(G)}\frac{1}{d_{i}(d_{i}-1)}\Big{)}=\frac{1}{n}\sum\limits_{i\in V(G)}\frac{\big{(}L(G)-1\big{)}(n-1)}{d_{i}(d_{i}-1)}.$
For pendant vertices we should add number of pendant $\frac{\text{number of pendant vertices}}{n}$ in the right side. Note that if there exist two vertices $i,j\in V(G)$ such that $d_{i}<d_{j}$ and $\operatorname{Str}(i)<\operatorname{Str}(j)$ then the inequality in this theorem will be strict. ∎
Let’s consider a star $G$ with $V(G)=n+1$ vertices. The star is geodetic graph. The central vertex has degree $n$, local clustering coefficient $c_{i}=0$ and stress centrality $\operatorname{Str}(i)=n(n-1)$. Other vertices are pendant ($d_{i}=1,\ c_{i}=0,\ \operatorname{Str}(i)=0$). Thus for this graph holds theorem 6.
$L(G)=\frac{n(2n-1)+n}{(n+1)n}=\frac{2n}{n+1},\quad C_{WS}=0.$
$1-C_{WS}(G)=1=\frac{\frac{n-1}{n+1}\,n}{n(n-1)}+\frac{n}{n+1}=\frac{1}{n+1}\sum\limits_{i\in V(G)}\frac{\big{(}L(G)-1\big{)}n}{d_{i}(d_{i}-1)}+\frac{\text{number of pendant vertices}}{n+1}.$
Thus, for this example holds equality.